OpenStudy (anonymous):

Fifty is what percent of 40?

OpenStudy (anonymous):

125%

OpenStudy (anonymous):

How did you get that? I dont want to cheat so can you tell me how you got it?

OpenStudy (anonymous):

divide 50 by 40

OpenStudy (anonymous):

Thanks mya

sammixboo (sammixboo):

OK so let's use this formula \(\rm \dfrac{part}{whole}=\dfrac{percent}{100}\) We know our part is 50 and or whole is 40 Plug in the values \(\rm \dfrac{50}{40}=\dfrac{percent}{100}\) Do you know how to solve this?

OpenStudy (anonymous):

yeah, divide

sammixboo (sammixboo):

Make sure not to give out direct answers and to guide the user to the answer isntead =)

sammixboo (sammixboo):

instead*

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

can you medal?

OpenStudy (anonymous):

No I dont@sammix i just kknow to do 50 divided by 40

sammixboo (sammixboo):

Yes that is all you need to do =)

OpenStudy (anonymous):

@Lucinda12345 stop spamming

OpenStudy (anonymous):

Ooh.I get it now.When you put the drawing up I looked at it and I got it!AAAAAAH!

OpenStudy (anonymous):

@myak598 is spamming

sammixboo (sammixboo):

=)